Solve the system of equations using the addition (elimination) method.
9x + 2y = 2
3x + 5y = 5
I am failing this class I have never had algebra before. please help

Answer by mukhopadhyay(490)

9x + 2y = 2 (multiply both sides by 5)
3x + 5y = 5 (multiply both sides by -2)
45x + 10y = 10
-6x - 10y = -10
By adding two rows above:
45x - 6x = 10-10
=>39x = 0
=> x = 0
Using first equation and substituting the value of x=0
9(0) + 2y = 2
=> y = 1
Answer x=0, y=1